# thanks: Corresponding authorthanks: Corresponding author

### v1

@@ II.1 The PAge model
where two free parameters $p_{\mathrm{age}}$ and $\eta$ are introduced to calibrate the expansion to satisfy Eq. (2). Their exact forms can be derived by combining the quadratic expansion of Eq. (1) with the constraint condition Eq. (2),
$p_{\mathrm{age}}=H_{0}t_{0},~{}\eta=1-\frac{3}{2}p^{2}_{\mathrm{age}}(1+q_{0}).$ (4)
The deceleration parameter $q_{0}=-1-\left.\dot{H}\,\right|_{t=t_{0}}/{H^{2}_{0}}$ appears in Eq. (4) arising from the quadratic term of Eq. (1). It is clear that $p_{\mathrm{age}}$ and $\eta$ quantify the extent to which the real universe deviates from the matter-dominated era relation $Ht=2/3$ in Eq. (3). To obtain $z(t)$, we can replace $H(t)$ in Eq. (3) with its definition $H(t)=-\mathrm{d}z/\mathrm{d}t/(1+z)$, and integrate both sides,
$1+z=\left(\frac{p_{\mathrm{age}}}{H_{0}t}\right)^{\frac{2}{3}}e^{\frac{1}{3}\left(1-\eta\frac{H_{0}t}{p_{\mathrm{age}}}\right)\left(3p_{\mathrm{age}}+\eta H_{0}t-\eta-2\right)}.$ (5)
Again, in Eq. (5), $p_{\mathrm{age}}$ and $\eta$ quantify the deviation from the matter-dominated universe $1+z=(t_{0}/{t})^{\frac{2}{3}}$.
The PAge model, as a type of Taylor expansion, can map specific dark energy models into a two-parameter space $(p_{\mathrm{age}},~{}\eta)$. Thus, the deviation from the $\Lambda\mathrm{CDM}$ model is clearly illustrated in the $(p_{\mathrm{age}},~{}\eta)$ parameter space. For simplicity, the mapping procedure is carried out by equating the deceleration parameter $q_{0}$ between the PAge model and the specific cosmological model, followed by computing $p_{\mathrm{age}}=H_{0}t_{0}$ for the reference model. Subsequently, $\eta$ can be derived from Eq. (4). More details about model mapping are provided in Ref. Luo et al. (2020).
In the following, we explicitly demonstrate how to map the specific dark energy models discussed in this paper into $(p_{\mathrm{age}},~{}\eta)$, including $\Lambda\mathrm{CDM}$, $w\mathrm{CDM}$, $w_{0}w_{a}\mathrm{CDM}$.
@@ II.2 Cosmological datasets
  • SNe: SNe are known as the standard candles due to their uniform intrinsic luminosity. The SNe data records the apparent magnitude $m_{\mathrm{B}}(z)$. The luminosity distance $D_{\mathrm{L}}(z)$ can be derived from the magnitude–redshift relation
    $m_{\mathrm{B}}(z)=5\mathrm{lg}\frac{D_{\mathrm{L}}(z)}{10~{}\mathrm{pc}}+M_{\mathrm{B}}=5\mathrm{lg}\frac{D_{\mathrm{L}}(z)H_{0}}{c}-5a_{\mathrm{B}},$ (10)
    where $M_{\mathrm{B}}$ is the absolute magnitude of SNe. $a_{\mathrm{B}}\equiv-\frac{1}{5}[\mathrm{lg}(c/10~{}\mathrm{pc}/H_{0})+M_{\mathrm{B}}]$ is the intercept of the magnitude–redshift relation Eq. (10). It is evident that $H_{0}$ and $M_{\mathrm{B}}$ are degenerate in $a_{\mathrm{B}}$, making it impossible to constrain both $H_{0}$ and $M_{\mathrm{B}}$ with SNe data alone. In IDL, we incorporate BAO and CC (described below) to break this degeneracy, avoiding the dependence on external prior.
@@
    We use two different SNe datasets, including PantheonPlus [^1] PantheonPlus data available at https://github.com/PantheonPlusSH0ES/DataRelease and DESY5222DESY5 data available at https://github.com/des-science/DES-SN5YR.
    The PantheonPlus SNe compilation consists of 1701 light curves of 1550 spectroscopically confirmed SNe Ia in the redshift range $2.260.001<z<$ 2.26 Scolnic et al. (2022). To reduce the effect of strong peculiar velocity dependence at low redshift Brout et al. (2022), only 1550 data points in the range $2.260.01<z<$ 2.26 are used.
    The DESY5 SNe sample is the largest and deepest single sample survey to date consisting of 1635 SNe Ia Abbott et al. (2024); Vincenzi et al. (2024); Sánchez et al. (2024) distributed in $1.30.10<z<$ 1.3 complemented by 194 low-redshift SNe Ia spanning $0.10.025<z<$ 0.1 Hicken et al. (2009, 2012); Krisciunas et al. (2017); Foley et al. (2018). The DESY5 data provide the Hubble diagram redshift, corrected apparent magnitudes $m^{\mathrm{corr.}}_{\mathrm{B}}(z)$, distance moduli $\mu(z)$ and the corresponding distance moduli errors $\sigma_{ij}$ and related quantities. Since the distance moduli contain no information about $M_{\mathrm{B}}$, we follow Refs. Huang et al. (2025b); Camilleri et al. (2025) to use the corrected apparent magnitudes to construct the distance moduli residual as $\Delta\mathbf{D}=m_{\mathrm{B}}^{\mathrm{corr.}}-M_{\mathrm{B}}-5\mathrm{lg}({D_{\mathrm{L}}(z)}/{10~{}\mathrm{pc}})$ to constrain $M_{\mathrm{B}}$.
  • BAO: The BAO records the imprint of early-universe plasma fluctuations in the late-time large-scale structure. BAO data provide three types of distances with respect to a fiducial cosmology, measured along different directions. Along the line-of-sight direction, BAO measures the Hubble distance
@@
    and along the transverse direction, BAO measures the (comoving) angular diameter distance
    $\frac{D_{\mathrm{M}}(z)}{r_{\mathrm{d}}}=\frac{(1+z)D_{\mathrm{A}}(z)}{r_{\mathrm{d}}}=\frac{D_{\mathrm{L}}}{(1+z)r_{\mathrm{d}}}.$ (13)
@@
    The SDSS BAO data are employed in this paper to cross-check our findings and connect with the existing literature. It includes two low-redshift data points 6dFGS Beutler et al. (2011), SDSS DR7 MGS Ross et al. (2015) and SDSS DR16 data Alam et al. (2021).
@@
    Tracer ${z_{\mathrm{eff}}}$ ${D_{\mathrm{M}}/r_{\mathrm{d}}}$ ${D_{\mathrm{H}}/r_{\mathrm{d}}}$ ${D_{\mathrm{V}}/r_{\mathrm{d}}}\text{ or }{r_{\mathrm{d}}/D_{\mathrm{V}}}$
@@
    BGS 0.30 - - 7.94 $\pm$ 0.08
    LRG1 0.51 13.59 $\pm$ 0.17 21.86 $\pm$ 0.43 -
    LRG2 0.71 17.35 $\pm$ 0.18 19.46 $\pm$ 0.33 -
    LRG3 + ELG1 0.93 21.58 $\pm$ 0.15 17.64 $\pm$ 0.19 -
    ELG2 1.32 27.60 $\pm$ 0.32 14.18 $\pm$ 0.22 -
    QSO 1.48 30.51 $\pm$ 0.76 12.82 $\pm$ 0.52 -
@@
    6dFGS 0.106 - - 0.336 $\pm$ 0.015
    MGS 0.15 - - 4.47 $\pm$ 0.17
    LRG 0.70 17.86 $\pm$ 0.33 19.33 $\pm$ 0.53 -
    ELG 0.85 - - $18.33^{+0.57}_{-0.62}$
    QS0 1.48 30.69 $\pm$ 0.80 13.26 $\pm$ 0.55 -
@@
    0.07 69 $\pm$ 19.6 Zhang et al. (2014)
    0.12 68.6 $\pm$ 26.2 Zhang et al. (2014)
    0.2 72.9 $\pm$ 29.6 Zhang et al. (2014)
    0.28 88.8 $\pm$ 36.6 Zhang et al. (2014)
@@
  • CC: CC measurements are incorporated to break the $r_{\mathrm{d}}$ and $H_{0}$ degeneracy in BAO data, serving a role analogous to the $r_{\mathrm{d}}$ prior in traditional IDL. It directly measures the Hubble parameter through massive passively evolving galaxies with old stellar populations and very low star formation rates Moresco (2023). By comparing two galaxies formed at the same time but separated by a small redshift interval, the Hubble parameter $H(z)$ can be calculated using the differential age method Jimenez and Loeb (2002), without assuming any cosmological model,
@@
    As CC data are not entirely independent, we incorporate correlations through the use of a covariance matrix when available. We adopt 32 CC data points, listed in Table 2, including 15 correlated measurements with the corresponding covariance matrix compiled in Ref. Moresco et al. (2020), and 17 uncorrelated ones. For the correlated measurements, we adopt the covariance matrix as described in Refs. Moresco et al. (2020, 2022),
@@
    $\chi_{\mathrm{CC}}^{2}=\sum_{i=1}^{17}\left[\frac{H_{\mathrm{obs}}\left(z_{i}\right)-H_{\mathrm{model}}\left(z_{i}\right)}{\sigma_{H,_{i}}}\right]^{2}+\Delta\mathbf{H^{T}}\mathrm{C}_{\mathrm{tot}}^{-1}\Delta\mathbf{H},$ (19)
@@
We consider four datasets, namely DESY5 + DESI DR2 + CC, DESY5 + SDSS + CC, PantheonPlus + DESI DR2 + CC, PantheonPlus + SDSS + CC. The final combined likelihood, $-2\mathrm{ln}\mathcal{L}=\chi^{2}_{\mathrm{SNe}}+\chi^{2}_{\mathrm{BAO}}+\chi^{2}_{\mathrm{CC}}$, is used to constrain the PAge model with five free parameters $\{\eta,~{}p_{\mathrm{age}},~{}M_{\mathrm{B}},~{}H_{0},~{}r_{\mathrm{d}}\}$. Then we employ the Markov Chain Monte Carlo (MCMC) sampler emcee [^5] https://github.com/dfm/emcee Foreman-Mackey et al. (2013) to explore the parameter space and obtain best-fit parameters through $\chi^{2}$ minimization. For comparison, we also consider the $\Lambda\mathrm{CDM}$ model, with four free parameters $\{\Omega_{\mathrm{m}},~{}M_{\mathrm{B}},~{}H_{0},~{}r_{\mathrm{d}}\}$. Uniform priors are adopted for the following parameter ranges $\eta\in[-2,2],~{}p_{\mathrm{age}}\in[0.15,2],~{}M_{\mathrm{B}}\in[-21,-19],~{}H_{0}\in[60,80]~{}\mathrm{km~{}s^{-1}~{}Mpc^{-1}},~{}r_{\mathrm{d}}\in[120,180]~{}\mathrm{Mpc},~{}\Omega_{\mathrm{m}}\in[0.2,0.4]$. Posterior samples are analyzed and visualized using GetDist [^6] https://github.com/cmbant/getdist Lewis (2019). For model comparison, we use the MCEvidence [^7] https://github.com/yabebalFantaye/MCEvidence algorithm as described in Ref. Heavens et al. (2017) to compute the Bayesian evidence directly from the MCMC chain.
Table 3: Cosmological constraints on the PAge and $\Lambda\mathrm{CDM}$ models based on the DESY5+DESI DR2+CC, DESY5+SDSS+CC, PantheonPlus+DESI DR2+CC, and PantheonPlus+SDSS+CC datasets. Here $H_{0}$ and $r_{\mathrm{d}}$ are in units of $\mathrm{km~{}s^{-1}~{}Mpc^{-1}}$ and $\mathrm{Mpc}$.
@@
DESY5+DESI DR2+CC 0.23 $\pm$ 0.04 0.94 $\pm$ 0.01 -19.36 $\pm$ 0.07 67.91 $\pm$ 2.33 145.05 $\pm$ 4.88
DESY5+SDSS+CC 0.16 $\pm$ 0.06 0.95 $\pm$ 0.01 -19.35 $\pm$ 0.07 67.74 $\pm$ 2.36 143.47 $\pm$ 4.99
PantheonPlus+DESI DR2+CC 0.31 $\pm$ 0.06 0.96 $\pm$ 0.01 -19.38 $\pm$ 0.07 68.97 $\pm$ 2.42 144.97 $\pm$ 4.89
PantheonPlus+SDSS+CC 0.25 $\pm$ 0.07 0.96 $\pm$ 0.02 -19.39 $\pm$ 0.08 68.58 $\pm$ 2.49 143.35 $\pm$ 4.99
@@
DESY5+DESI DR2+CC 0.32 $\pm$ 0.01 -19.35 $\pm$ 0.07 68.99 $\pm$ 2.35 145.18 $\pm$ 4.89
DESY5+SDSS+CC 0.35 $\pm$ 0.01 -19.38 $\pm$ 0.07 67.46 $\pm$ 2.37 143.42 $\pm$ 4.99
PantheonPlus+DESI DR2+CC 0.30 $\pm$ 0.01 -19.37 $\pm$ 0.07 69.75 $\pm$ 2.39 144.98 $\pm$ 4.89
PantheonPlus+SDSS+CC 0.32 $\pm$ 0.02 -19.39 $\pm$ 0.08 69.08 $\pm$ 2.51 143.16 $\pm$ 4.99

### v2

@@ II.1 The PAge and MAPAge parametrization
where $\eta$ quantifies the extent to which the real universe deviates from the pure matter universe with $\eta=0$ in Eq. (7). To obtain $z(t)$, we can replace $H(t)$ in Eq. (7) with its definition $H(t)=-\mathrm{d}z/\mathrm{d}t/(1+z)$, and integrate both sides,
$1+z=\left(\frac{p_{\mathrm{age}}}{H_{0}t}\right)^{2/3}\mathrm{e}^{\frac{1}{3}\left(1-\frac{H_{0}t}{p_{\mathrm{age}}}\right)\left(3p_{\mathrm{age}}+\eta\frac{H_{0}t}{p_{\mathrm{age}}}-\eta-2\right)}.$ (8)
Again, in Eq. (8), $\eta$ quantify the deviation from the matter-dominated Universe $1+z=(t_{0}/{t})^{\frac{2}{3}}$.
Similarly, we can parametrize $H(t)t$ with cubic function of $t$, which is referred to as MAPAge in Ref. Huang et al. (2021a). It takes the following forms:
$\frac{H}{H_{0}}=1+\frac{2}{3}$ $\left[1-(\eta_{1}+\eta_{2})\frac{H_{0}t}{p_{\mathrm{age}}}+\eta_{2}\left(\frac{H_{0}t}{p_{\mathrm{age}}}\right)^{2}\right]$
$\times\left(\frac{1}{H_{0}t}-\frac{1}{p_{\mathrm{age}}}\right),$ (9)
where $\eta_{1}$ equals to $\eta$ in PAge. $\eta_{2}$ is introduced to further parametrize $H(t)t$ and is given by
$\eta_{2}=1-\frac{3}{4}p^{2}_{\mathrm{age}}(2+j_{0}+3q_{0}),$ (10)
Figure 1: Comparison of approximation accuracy for PAge (green dash-dotted), MAPAge (red dashed), Taylor expansion in $z$ (brown dotted), and $y$ (light beige), targeting the fiducial $w_{0}w_{a}\mathrm{CDM}$ (black) model. The lower-right inset figure shows normalized Hubble parameter $H(z)/H(z)_{w_{0}w_{a}\mathrm{CDM}}$ up to $z\sim 10^{3}$. We fix the parameters of the $w_{0}w_{a}\rm CDM$ model to DESI DR2 results: $\Omega_{\mathrm{m}}=0.32,~H_{0}=66.74~\mathrm{km~s^{-1}~Mpc^{-1}},~w_{0}=-0.75,~w_{a}=-0.86$ Abdul Karim et al. (2025).
The PAge model can map specific dark energy models into a two-parameter space $(p_{\mathrm{age}},~\eta)$. Thus, the deviation from the $\Lambda\mathrm{CDM}$ model is clearly illustrated in the $(p_{\mathrm{age}},~\eta)$ parameter space. For simplicity, the mapping procedure is carried out by equating the deceleration parameter $q_{0}$ between the PAge model and the specific cosmological model, followed by computing $p_{\mathrm{age}}=H_{0}t_{0}$ for the reference model. Subsequently, $\eta$ can be derived from Eq. (6). More details about model mapping are provided in Ref. Luo et al. (2020).
In the following, we explicitly demonstrate how to map the specific dark energy models discussed in this paper into $(p_{\mathrm{age}},~\eta)$, including $\Lambda\mathrm{CDM}$, $w\mathrm{CDM}$, and $w_{0}w_{a}\mathrm{CDM}$.
@@ II.2 Cosmological datasets
II.2.1 SNe
SNe are known as the standard candles due to their uniform intrinsic luminosity. The SNe data records the apparent magnitude $m_{\mathrm{B}}(z)$. The luminosity distance $D_{\mathrm{L}}(z)$ can be derived from the magnitude–redshift relation
$m_{\mathrm{B}}(z)=5~\mathrm{lg}\frac{D_{\mathrm{L}}(z)}{10~\mathrm{pc}}+M_{\mathrm{B}}=5~\mathrm{lg}\frac{D_{\mathrm{L}}(z)H_{0}}{c}-5a_{\mathrm{B}},$ (15)
where $M_{\mathrm{B}}$ is the absolute magnitude of SNe. $a_{\mathrm{B}}\equiv-\frac{1}{5}[5~\mathrm{lg}(c/10~\mathrm{pc}/H_{0})+M_{\mathrm{B}}]$ is the intercept of the magnitude–redshift relation Eq. (15). It is evident that $H_{0}$ and $M_{\mathrm{B}}$ are degenerate in $a_{\mathrm{B}}$, making it impossible to constrain both $H_{0}$ and $M_{\mathrm{B}}$ with SNe data alone. In IDL, we incorporate BAO and CC (described below) to break this degeneracy, avoiding the dependence on external prior.
@@ II.2.1 SNe
We use two different SNe datasets, including PantheonPlus [^3] PantheonPlus data available at https://github.com/PantheonPlusSH0ES/DataRelease and DESY5444DESY5 data available at https://github.com/des-science/DES-SN5YR.
The PantheonPlus SNe compilation consists of 1701 light curves of 1550 spectroscopically confirmed SNe Ia in the redshift range $0.001<z<2.26$ Scolnic et al. (2022). To reduce the effect of strong peculiar velocity dependence at low redshift Brout et al. (2022), only 1550 data points in the range $0.01<z<2.26$ are used.
The DESY5 SNe sample is the largest and deepest single sample survey to date consisting of 1635 SNe Ia Abbott et al. (2024); Vincenzi et al. (2024); Sánchez et al. (2024) distributed in $0.10<z<1.3$ complemented by 194 low-redshift SNe Ia spanning $0.025<z<0.1$ Hicken et al. (2009, 2012); Krisciunas et al. (2017); Foley et al. (2018). The DESY5 data provide the Hubble diagram redshift, corrected apparent magnitudes $m^{\mathrm{corr.}}_{\mathrm{B}}(z)$, distance moduli $\mu(z)$, and the corresponding distance moduli errors $\sigma_{ij}$ and related quantities. Since the distance moduli contain no information about $M_{\mathrm{B}}$, we follow Refs. Huang et al. (2025b); Camilleri et al. (2025) to use the corrected apparent magnitudes to construct the distance moduli residual as $\Delta\mathbf{D}=m_{\mathrm{B}}^{\mathrm{corr.}}-M_{\mathrm{B}}-5\mathrm{lg}({D_{\mathrm{L}}(z)}/{10~\mathrm{pc}})$ to constrain $M_{\mathrm{B}}$.
II.2.2 BAO
The BAO records the imprint of early-Universe plasma fluctuations in the late-time large-scale structure. BAO data provide three types of distances with respect to a fiducial cosmology, measured along different directions. Along the line-of-sight direction, BAO measures the Hubble distance,
@@ II.2.2 BAO
and along the transverse direction, BAO measures the (comoving) angular diameter distance,
$\frac{D_{\mathrm{M}}(z)}{r_{\mathrm{d}}}=\frac{(1+z)D_{\mathrm{A}}(z)}{r_{\mathrm{d}}}=\frac{D_{\mathrm{L}}(z)}{(1+z)r_{\mathrm{d}}}.$ (18)
@@
The SDSS BAO data are employed in this paper to cross-check our findings and connect with the existing literature. It includes two low-redshift data points 6dFGS Beutler et al. (2011), SDSS DR7 MGS Ross et al. (2015), and SDSS DR16 data Alam et al. (2021).
@@
[…]
@@
BGS 0.30 $\cdots$ $\cdots$ 7.94 $\pm$ 0.08
LRG1 0.51 13.59 $\pm$ 0.17 21.86 $\pm$ 0.43 $\cdots$
LRG2 0.71 17.35 $\pm$ 0.18 19.46 $\pm$ 0.33 $\cdots$
LRG3 + ELG1 0.93 21.58 $\pm$ 0.15 17.64 $\pm$ 0.19 $\cdots$
ELG2 1.32 27.60 $\pm$ 0.32 14.18 $\pm$ 0.22 $\cdots$
QSO 1.48 30.51 $\pm$ 0.76 12.82 $\pm$ 0.52 $\cdots$
@@
6dFGS 0.106 $\cdots$ $\cdots$ 0.336 $\pm$ 0.015
MGS 0.15 $\cdots$ $\cdots$ 4.47 $\pm$ 0.17
LRG 0.70 17.86 $\pm$ 0.33 19.33 $\pm$ 0.53 $\cdots$
ELG 0.85 $\cdots$ $\cdots$ $18.33^{+0.57}_{-0.62}$
QS0 1.48 30.69 $\pm$ 0.80 13.26 $\pm$ 0.55 $\cdots$
@@
0.07 69 $\pm$ 19.6 Zhang et al. (2014b)
0.12 68.6 $\pm$ 26.2 Zhang et al. (2014b)
0.2 72.9 $\pm$ 29.6 Zhang et al. (2014b)
0.28 88.8 $\pm$ 36.6 Zhang et al. (2014b)
@@
II.2.3 CC
CC measurements are incorporated to break the $r_{\mathrm{d}}$ and $H_{0}$ degeneracy in BAO data, serving a role analogous to the $r_{\mathrm{d}}$ prior in traditional IDL. It directly measures the Hubble parameter through massive passively evolving galaxies with old stellar populations and very low star formation rates Moresco (2023). By comparing two galaxies formed at the same time but separated by a small redshift interval, the Hubble parameter $H(z)$ can be calculated using the differential age method Jimenez and Loeb (2002), without assuming any cosmological model,
@@
As CC data are not entirely independent, we incorporate correlations through the use of a covariance matrix when available. We adopt 32 CC data points, listed in Table 2, including 15 correlated measurements with the corresponding covariance matrix compiled in Ref. Moresco et al. (2020) and 17 uncorrelated ones. For the correlated measurements, we adopt the covariance matrix as described in Refs. Moresco et al. (2020, 2022),
@@
$\chi_{\mathrm{CC}}^{2}=\sum_{i=1}^{17}\left[\frac{H_{\mathrm{obs}}\left(z_{i}\right)-H_{\mathrm{model}}\left(z_{i}\right)}{\sigma_{H,_{i}}}\right]^{2}+\Delta\mathbf{H}^{\rm T}\mathrm{C}_{\mathrm{tot}}^{-1}\Delta\mathbf{H},$ (24)
@@
We consider four datasets, namely DESY5 + DESI DR2 + CC, DESY5 + SDSS + CC, PantheonPlus + DESI DR2 + CC, and PantheonPlus + SDSS + CC. The final combined likelihood, $2\mathrm{ln}\mathcal{L}=\chi^{2}_{\mathrm{SNe}}+\chi^{2}_{\mathrm{BAO}}+\chi^{2}_{\mathrm{CC}}$, is used to constrain the PAge model with five free parameters $\{\eta,~p_{\mathrm{age}},~M_{\mathrm{B}},~H_{0},~r_{\mathrm{d}}\}$. Then we employ the Markov Chain Monte Carlo (MCMC) sampler emcee [^7] https://github.com/dfm/emcee Foreman-Mackey et al. (2013) to explore the parameter space and obtain best-fit parameters through $\chi^{2}$ minimization. For comparison, we also consider the $\Lambda\mathrm{CDM}$ model, with four free parameters $\{\Omega_{\mathrm{m}},~M_{\mathrm{B}},~H_{0},~r_{\mathrm{d}}\}$. Uniform priors are adopted for the following parameter ranges $\eta\in[-2,2],~p_{\mathrm{age}}\in[0.15,2],~M_{\mathrm{B}}\in[-21,-19],~H_{0}\in[60,80]~\mathrm{km~s^{-1}~Mpc^{-1}},~r_{\mathrm{d}}\in[120,180]~\mathrm{Mpc},~\Omega_{\mathrm{m}}\in[0.2,0.4]$. Posterior samples are analyzed and visualized using GetDist [^8] https://github.com/cmbant/getdist Lewis (2025). For model comparison, we use the MCEvidence [^9] https://github.com/yabebalFantaye/MCEvidence algorithm as described in Ref. Heavens et al. (2017) to compute the Bayesian evidence directly from the MCMC chain.
Table 3: Cosmological constraints on the PAge and $\Lambda\mathrm{CDM}$ models based on the DESY5+DESI DR2+CC, DESY5+SDSS+CC, PantheonPlus+DESI DR2+CC, and PantheonPlus+SDSS+CC datasets. Here, $H_{0}$ and $r_{\mathrm{d}}$ are in units of $\mathrm{km~s^{-1}~Mpc^{-1}}$ and $\mathrm{Mpc}$.
@@
DESY5+DESI DR2+CC 0.23 $\pm$ 0.04 0.94 $\pm$ 0.01 $\cdots$ $19.36\pm$ 0.07 67.91 $\pm$ 2.33 $145.0^{+4.5}_{-5.2}$
DESY5+SDSS+$C$ 0.16 $\pm$ 0.06 0.95 $\pm$ 0.01 $\cdots$ $19.35\pm$ 0.07 67.74 $\pm$ 2.36 $143.5^{+4.6}_{-5.3}$
PantheonPlus+DESI DR2+CC 0.31 $\pm$ 0.06 0.96 $\pm$ 0.01 $\cdots$ $19.38\pm$ 0.07 68.97 $\pm$ 2.42 $145.0^{+4.5}_{-5.2}$
PantheonPlus+SDSS+$C$ 0.25 $\pm$ 0.07 0.96 $\pm$ 0.02 $\cdots$ $19.39\pm$ 0.08 68.58 $\pm$ 2.49 $143.4^{+4.6}_{-5.3}$
@@
DESY5+DESI DR2+CC $\cdots$ $\cdots$ 0.32 $\pm$ 0.01 $19.35\pm$ 0.07 68.99 $\pm$ 2.35 $145.2^{+4.5}_{-5.1}$
DESY5+SDSS+$C$ $\cdots$ $\cdots$ 0.35 $\pm$ 0.01 $19.38\pm$ 0.07 67.46 $\pm$ 2.37 $143.4^{+4.6}_{-5.3}$
PantheonPlus+DESI DR2+CC $\cdots$ $\cdots$ 0.30 $\pm$ 0.01 $19.37\pm$ 0.07 69.75 $\pm$ 2.39 $145.0^{+4.5}_{-5.2}$
PantheonPlus+SDSS+$C$ $\cdots$ $\cdots$ 0.32 $\pm$ 0.02 $19.39\pm$ 0.08 69.08 $\pm$ 2.51 $143.2^{+4.6}_{-5.3}$